### v1

# Parallelizing the Approximate Minimum Degree Ordering Algorithm: Strategies and Evaluation

Abstract.
The approximate minimum degree algorithm is widely used before numerical factorization to reduce fill-in for sparse matrices. While considerable attention has been given to the numerical factorization process, less focus has been placed on parallelizing the approximate minimum degree algorithm itself. In this paper, we explore different parallelization strategies, and introduce a novel parallel framework that leverages multiple elimination on distance-2 independent sets. Our evaluation shows that parallelism within individual elimination steps is limited due to low computational workload and significant memory contention. In contrast, our proposed framework overcomes these challenges by parallelizing the work across elimination steps. To the best of our knowledge, our implementation is the first scalable shared memory implementation of the approximate minimum degree algorithm. Experimental results show that we achieve up to an 8.30x speedup using 64 threads over the state-of-the-art sequential implementation in SuiteSparse.
1. Introduction
Efficiently solving sparse linear systems is a cornerstone of many scientific and engineering applications. Cholesky factorization, a widely used direct method for solving symmetric positive definite systems, often encounters challenges when the given matrix $A$ is sparse. In such cases, zero entries in $A$ can transform into nonzeros, known as fill-ins, in the Cholesky factor. These fill-ins can significantly increase both memory usage and computational cost, undermining the efficiency of the factorization. To address this, fill-reducing orderings seek to find a permutation matrix $P$ such that the permuted matrix $PAP^{T}=LL^{T}$ reduces the number of fill-ins in the Cholesky factor $L$. However, minimizing the number of fill-ins is an NP-complete problem (Yannakakis, 1981), and as a result, various heuristics have been developed, such as nested dissection (George, 1973; Gilbert and Tarjan, 1986) and approximate minimum degree (AMD) (Amestoy et al., 1996). This paper solely focuses on parallelizing the AMD algorithm in shared memory since to the best of our knowledge, no parallel implementation currently exists.
The need for a scalable shared memory parallel AMD implementation is motivated by literature and practical use cases. Trotter et al. (Trotter et al., 2023) recently conducted an extensive study on 490 matrices and found that while nested dissection and the AMD algorithm achieved similar levels of fill reduction, the ordering time for nested dissection was up to an order of magnitude longer than that of AMD. Furthermore, while significant work has been done to parallelize nested dissection on both shared and distributed memory systems (Chevalier and Pellegrini, 2008; Karypis and Kumar, 1998; LaSalle and Karypis, 2015), no similar efforts have been made for the AMD algorithm, highlighting a key gap in the literature. Meanwhile, GPU Cholesky solvers have steadily improved over time, as shown in Table 1, driven by advances in software, and are expected to continue accelerating with future GPU generations. In contrast, the standalone sequential AMD implementation on CPU has emerged as a critical bottleneck. Our goal in this paper is therefore to design a scalable shared memory parallel AMD implementation, closing the gap in literature and in practice.
Table 1. Time (in seconds) for the sequential AMD implementation from SuiteSparse (v7.10.1) (Davis, 2019) on an AMD EPYC 7763 CPU, and for Cholesky solver cuSolverSp (Corporation, 2025b) (v11.6.1) and its successor cuDSS (Corporation, 2025a) (v0.5.0) to solve the reordered symmetric positive definite system on an A100 GPU. The matrices are randomly permuted symmetrically beforehand to decouple tie-breaking issues in the AMD algorithm.
Matrix Name SuiteSparse AMD cuSolverSp cuDSS
nd24k 0.83 117.17 4.11
ldoor 1.28 11.01 3.55
Flan_1565 4.18 out of memory 16.85
Parallelizing the AMD algorithm presents two main challenges. First, each elimination step involves selecting a pivot vertex with the minimum approximate degree, eliminating it, and then updating its neighborhood and the approximate degrees of the neighbors. These steps are inherently sequential, as a new pivot cannot be selected until all updates from the previous pivot are complete. While neighborhood and approximate degree updates could, in theory, be parallelized, the opportunities for efficient parallelism are limited due to little amount of work and high memory contention. Second, although the multiple elimination strategy from the sequential multiple minimum degree algorithm (Liu, 1985) can help reduce dependencies by selecting a maximal independent set of pivots, this approach introduces significant overlap in these pivots’ neighborhoods. While such overlap is beneficial in the sequential setting—reducing the cost of degree updates—it introduces contention on the underlying graph structure and complicates the maintenance of approximate degrees, ultimately hindering scalability when attempting to parallelize the AMD algorithm using multiple elimination.
In this paper, we address these challenges by selecting distance-2 independent sets of pivots, enabling parallel elimination steps while keeping approximate degree updates simple and efficient. We further exploit key properties of the AMD algorithm to design specialized concurrent data structures—including a concurrent graph representation and concurrent approximate degree lists—that eliminate the need for dynamic memory allocation and significantly reduce memory contention. Together, these optimizations lead to the first scalable shared memory implementation of the AMD algorithm, outperforming the long-standing sequential version in SuiteSparse (Davis, 2019). Our contributions include:
  • We present the first scalable shared memory implementation of the AMD algorithm. Our approach leverages multiple elimination on distance-2 independent sets to extract parallelism across elimination steps, and introduces specialized concurrent data structures tailored to the unique properties of the AMD algorithm.
  • Our implementation achieves ordering quality comparable to the state-of-the-art, with tunable control via relaxation and limitation factors to balance fill-in and parallelism.
  • We demonstrate up to an 8.30×8.30$\times$8.30 × speedup using 64 threads on a diverse set of real-world sparse matrices compared to the leading sequential implementation in SuiteSparse. Our in-depth performance analysis highlights current bottlenecks and points to promising directions for future research.
2. Preliminaries
Before introducing the approximate minimum degree (AMD) algorithm (Amestoy et al., 1996), we begin with its predecessors, the minimum degree algorithm (Rose, 1972) and the multiple minimum degree algorithm (Liu, 1985), to establish the foundational concepts. From this point onward, we disregard the positive definiteness of the matrices, as it does not impact the ordering algorithms discussed in this paper.
Figure 1. An example illustrating how elimination graphs work. For demonstration purposes, we eliminate vertices 5, 2, and 9 in order, rather than following the minimum degree criterion. When a vertex is eliminated, its neighbors form a clique.
[…]
An example to demonstrate how elimination graphs work.
Figure 2. An example illustrating how quotient graphs work. For demonstration purposes, we eliminate variables 5, 2, and 9 in order, rather than following the minimum degree criterion. Circles represent variables and squares represent elements.
[…]
An example to demonstrate how quotient graphs work.
2.1. Minimum Degree Algorithm
The minimum degree algorithm is most naturally described using elimination graphs (Rose, 1972). Let $A$ be an $n\times n$ symmetric sparse matrix. The sparsity pattern of $A$ can be represented by a graph $G^{0}=(V^{0},E^{0})$, where $V^{0}=\{1,2,\cdots,n\}$, and each vertex $i$ corresponds to row/column $i$ of $A$. An edge $(i,j)$ is in $E^{0}$ if and only if $A_{i,j}\neq 0$. The elimination graph $G^{k}=(V^{k},E^{k})$ represents the sparsity pattern of the remaining submatrix after the first $k$ pivots have been selected and eliminated during Cholesky factorization. Let $N^{k}_{v}$ denote the neighborhood of vertex $v$ in the elimination graph $G^{k}$. The elimination process begins with $G^{0}$. To obtain $G^{k}$ from $G^{k-1}$, the $k$th pivot $p$ is selected from $V^{k-1}$, removed from the vertex set, and its neighbors are connected to form a clique—reflecting the fill-ins introduced by an outer product update in Cholesky factorization. An example is presented in Figure 1. Formally, the update can be expressed as:
$V^{k}$ $=V^{k-1}\setminus\{p\},$
$E^{k}$ $=\left(V^{k}\times V^{k}\right)\cap\left(E^{k-1}\cup\left(N^{k-1}_{p}\times N^{k-1}_{p}\right)\right).$
The minimum degree algorithm selects, at each step, the vertex with the smallest degree in $G^{k-1}$ as the $k$th pivot. Since multiple vertices can have the same degree, a tie-breaking strategy is required. Ties are typically resolved arbitrarily or based on the initial ordering; however, it has been shown that the choice of tie-breaking strategy can significantly affect the number of fill-ins produced (George and Liu, 1989).
2.2. Quotient Graphs
A key challenge with elimination graphs is that the storage needed for $G^{k}$ is not known in advance due to the formation of cliques during elimination, which necessitates dynamic memory management. The quotient graph (George and Liu, 1980), also known as the generalized element model or superelement (Eisenstat et al., 1976; Speelpenning, 1978; Duff and Reid, 1983), addresses this issue by representing cliques more compactly—storing only the set of vertices in each clique. George and Liu (George and Liu, 1981) showed that the quotient graph takes no more storage than the original graph $G^{0}$.
Following historical terminology, we refer to eliminated vertices from the original graph as elements, and those yet to be eliminated as variables. To stay consistent with the literature, we adopt the notations used in (Amestoy et al., 1996). The quotient graph $\mathcal{G}^{k}=\left(V^{k},\overline{V}^{k},\mathcal{A}^{k},\mathcal{E}^{k},\mathcal{L}^{k}\right)$ provides an implicit representation of the elimination graph $G^{k}$. Here, $V^{k}$ denotes the set of variables, $\overline{V}^{k}$ the set of elements, $\mathcal{A}^{k}$ the set of variable-to-variable connections, $\mathcal{E}^{k}$ the connections from variables to elements, and $\mathcal{L}^{k}$ the connections from elements to variables. Notably, there is no connection between elements, and while $\mathcal{E}^{k}$ and $\mathcal{L}^{k}$ are essentially transposes of each other, they are often maintained separately in practical implementations for efficiency. In essence, the variables $V^{k}$ correspond to the uneliminated vertices, with $\mathcal{A}^{k}$ being the remaining edges from the original graph; the elements $\overline{V}^{k}$ represent cliques formed during elimination, with their connections compactly encoded in $\mathcal{E}^{k}$ and $\mathcal{L}^{k}$.
The elimination process begins with the initial quotient graph $\mathcal{G}^{0}=(V^{0},\emptyset,E^{0},\emptyset,\emptyset)$. For simplicity, we omit the superscript $k$ when the context is clear. Given a variable $v$ and an element $e$ in the quotient graph, we define the following adjacency sets:
$\mathcal{A}_{v}$ =$=\{v^{\prime}\mid(v,v^{\prime})\in\mathcal{A}\}\subseteq V,$
$\mathcal{E}_{v}$ =$=\{e^{\prime}\mid(v,e^{\prime})\in\mathcal{E}\}\subseteq\overline{V},$
$\mathcal{L}_{e}$ =$=\{v^{\prime}\mid(e,v^{\prime})\in\mathcal{L}\}\subseteq V.$
The quotient graph and the elimination graph are closely related through the following invariant: for any vertex $v$ in $G$, which also appears as a variable in $\mathcal{G}$, its neighborhood in the elimination graph can be recovered by
(1) $N_{v}=\left(\mathcal{A}_{v}\cup\bigcup_{e\in\mathcal{E}_{v}}\mathcal{L}_{e}\right)\setminus\{v\}.$
In other words, the neighborhood of $v$ in the elimination graph is reconstructed from the quotient graph by interpreting the remaining original edges and the clique information encoded in the adjacent elements.
When a variable $p$ is selected as the $k$th pivot from $\mathcal{G}^{k-1}$, it becomes an element in $\mathcal{G}^{k}$. Its neighborhood $N_{p}$ is computed using Equation (1). To preserve the invariant for the remaining variables, the clique induced by $p$, namely $\{p\}\times N_{p}$, is added to $\mathcal{L}$, and similarly $N_{p}\times\{p\}$ is added to $\mathcal{E}$. Only the set of variables forming the clique is stored, instead of their pairwise connections. Additionally, connections now covered by this newly formed clique are removed. Formally, the connection updates are defined as:
$\mathcal{A}^{k}$ $=\left(\mathcal{A}^{k-1}\setminus\left(N_{p}^{k-1}\times N_{p}^{k-1}\right)\right)\cap\left(V^{k}\times V^{k}\right),$
$\mathcal{E}^{k}$ $=\left(\mathcal{E}^{k-1}\setminus\bigcup_{e\in\mathcal{E}_{p}^{k-1}}\left(\mathcal{L}^{k-1}_{e}\times\{e\}\right)\right)\cup\left(N^{k-1}_{p}\times\{p\}\right),$
$\mathcal{L}^{k}$ $=\left(\mathcal{L}^{k-1}\setminus\bigcup_{e\in\mathcal{E}_{p}^{k-1}}\left(\{e\}\times\mathcal{L}^{k-1}_{e}\right)\right)\cup\left(\{p\}\times N^{k-1}_{p}\right).$
Figure 2 illustrates how quotient graphs work, where circles represent variables and squares represent elements. Taking $\mathcal{G}^{1}$ as an example, we determine $N_{2}=\{1,3,4,6,8\}$ based on Equation (1) using the information encoded in element 5 and the original edges. This result is consistent with $G^{1}$ shown in Figure 1. Once variable 2 is eliminated, a new adjacency set $\mathcal{L}_{2}=N_{2}$ is formed. Redundant variable-variable connections, such as the connection between 1 and 4, are removed in the new quotient graph $\mathcal{G}^{2}$, as they can be implied from $\mathcal{L}_{2}$. Additionally, element 5 is removed since the clique it encoded is now covered by the new clique represented by element 2 using $\mathcal{L}_{2}$.
To build the minimum degree algorithm using quotient graphs, it is necessary to update the degrees of all variables $v\in N_{p}$ after eliminating the pivot $p$. However, this step is computationally expensive because reconstructing the neighborhood of each $v$ requires performing set union operations, as described in Equation (1). The cost of these operations is linear in the total size of the sets involved. These degree updates have been identified as the primary bottleneck of the minimum degree algorithm in the literature (Liu, 1985; George and Liu, 1989).
2.3. Multiple Minimum Degree Algorithm
The multiple minimum degree algorithm (Liu, 1985) addresses the bottleneck in degree updates by selecting a maximal independent set of pivots, denoted by $M$, within an additive relaxation of the minimum degree criterion in the elimination graph. All pivots in $M$ are eliminated sequentially with their associated connection updates applied before any degree updates are performed on affected variables. The independence of these pivots ensures that pivots do not interfere with one another, while the maximality of the set increases the likelihood of overlapping pivots’ neighborhoods. This overlap is key to reducing the cost of degree updates: whereas the standard minimum degree algorithm requires a separate degree update for each adjacent pivot in $M$, the multiple minimum degree algorithm defers degree updates until all pivots in $M$ are eliminated, allowing each affected variable to perform only a single, consolidated degree update.
For example, consider selecting a maximal independent set $M=\{1,3,7,9\}$ from $\mathcal{G}^{0}$ in Figure 2. If the variables in $M$ are eliminated one by one using the minimum degree algorithm, then each of the variables 2, 4, 6, and 8 will incur two degree updates. In contrast, with multiple elimination, where all variables in $M$ are eliminated in a single step, only one degree update is required for each of the variables 2, 4, 6, and 8 after all connection updates are applied.
Figure 3. An example illustrating how to compute approximate degrees after eliminating variable 7. To determine the exact degree of variable 8 in $\mathcal{G}^{4}$, which is 4, one would need to compute $|\mathcal{L}_{2}\cup\mathcal{L}_{7}\cup\mathcal{L}_{9}\setminus\{8\}|$. A naïve estimate using a union bound results in double counting, as variables 4 and 6 appear twice, yielding an estimate of 6. The approximate degree, however, mitigates such double counting by leveraging information already captured in $\mathcal{L}_{7}$ (from the pivot 7). Specifically, the estimate is computed as $|\mathcal{L}_{7}\setminus\{8\}|+|\mathcal{L}_{2}\setminus\mathcal{L}_{7}|+|\mathcal{L}_{9}\setminus\mathcal{L}_{7}|=5$, with only variable 6 being double counted.
[…]
An example to demonstrate how approximate minimum degrees work.
2.4. Approximate Minimum Degree Algorithm
The approximate minimum degree (AMD) algorithm (Amestoy et al., 1996) adopts an alternative strategy to the multiple minimum degree algorithm for reducing the cost of degree updates. Instead of computing exact degrees for all variables $v$ in $\mathcal{L}_{p}=N^{k-1}_{p}$ after the connection updates, the AMD algorithm estimates an upper bound, $d^{k}_{v}$, for the exact degree:
$d^{k}_{v}=\min\left\{\begin{array}[]{l}n-k-1,\\ d^{k-1}_{v}+|\mathcal{L}_{p}\setminus\{v\}|-1,\\ |\mathcal{A}_{v}\setminus\{v\}|+|\mathcal{L}_{p}\setminus\{v\}|+\sum_{e\in\mathcal{E}_{v}}|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|\end{array}\right\}$
with $d^{0}_{v}$ initialized as the exact degree of $v$ in the original graph. Note that $d_{v}$ remains unchanged if $v\notin\mathcal{L}_{p}$. This estimate is known as the approximate degree. The first term reflects the maximum possible degree given the size of the remaining submatrix, the second accounts for the worst case fill-in introduced by $p$, and the third refines the estimate by avoiding double counting using local information from the neighborhood of $p$. An example is shown in Figure 3.
To efficiently compute $|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|$, Algorithm 1 uses the identity $|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|=|\mathcal{L}_{e}|-|\mathcal{L}_{e}\cap\mathcal{L}_{p}|$. During the scan, if an element $e$ is encountered for the first time, $w(e)$ is initialized to $|\mathcal{L}_{e}|$, and then decremented once for every variable $v\in\mathcal{L}_{e}\cap\mathcal{L}_{p}$. Otherwise, $w(e)$ remains negative. Combining both cases, the final value of $|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|$ for each $e\in\overline{V}$ is given by (Amestoy et al., 1996):
$|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|=\left\{\begin{array}[]{ll}w(e)&\text{if }w(e)\geq 0,\\ |\mathcal{L}_{e}|&\text{otherwise}\end{array}\right\}.$
Algorithm 1 Computation of $|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|$ for all $e\in\overline{V}$ (Amestoy et al., 1996)
[…]
2:$w(e)\leftarrow|\mathcal{L}_{e}|$EndIf
[…]
By maintaining approximate degrees, the AMD algorithm selects, at each step, the variable with the smallest approximate degree in $\mathcal{G}^{k-1}$ as the $k$th pivot. This approach significantly reduces the cost of degree updates, as it avoids explicitly accessing the adjacency sets $\mathcal{L}_{e}$, which are required for exact degree computation in Equation (1) for the minimum degree algorithm. For a detailed complexity analysis of both the AMD and the minimum degree algorithms, we refer the reader to (Heggernes et al., 2002).
The AMD algorithm also incorporates several additional techniques, including mass elimination (George and McIntyre, 1978), element absorption (Duff and Reid, 1983), indistinguishable nodes (George and Liu, 1981), and external degrees (Liu, 1985). These techniques are included in our parallel implementation of the AMD algorithm, though we omit detailed discussion here, as they are orthogonal to parallelization and may be explored in the cited references for interested readers.
2.5. Related Work
2.5.1. Minimum Degree and its Variants
The minimum degree heuristic, a symmetric analog to Markowitz’s algorithm (Markowitz, 1957), was first proposed by Tinney and Walker (Tinney and Walker, 1967). A graph-theoretical interpretation was later developed by Rose (Rose, 1972), who coined the term minimum degree algorithm, as it selected the pivot with the smallest degree at each elimination step. Over the years, numerous enhancements have been proposed and have become standard components in modern implementations. These include mass elimination (George and McIntyre, 1978), indistinguishable nodes (George and Liu, 1981), quotient graphs (George and Liu, 1980) (also known as generalized element models or superelements (Eisenstat et al., 1976; Speelpenning, 1978; Duff and Reid, 1983)), incomplete degree updates (Liu, 1985; Eisenstat et al., 1982), element absorption (Duff and Reid, 1983), multiple elimination (Liu, 1985), and external degrees (Liu, 1985). These techniques and their impacts were comprehensively summarized by George and Liu (George and Liu, 1989).
2.5.2. Approximate Minimum Degree and its Variants
Since the degree updates have been identified as the primary bottleneck of the minimum degree algorithm in the literature (Liu, 1985; George and Liu, 1989), several researchers (Davis and Duff, 1997, 1991; Gilbert et al., 1992; Amestoy et al., 1996) proposed computing upper bounds on the degrees rather than computing exact values. Among these approaches, the AMD algorithm (Amestoy et al., 1996) has been shown to be the most effective. Its sequential implementation in SuiteSparse (Davis, 2019) is now widely used in modern sparse solvers. The success of the AMD algorithm also spurred further interest in nonsymmetric matrices, leading to the development of the column approximate minimum degree algorithm (Davis et al., 2004), which aims to reorder columns to minimize worst-case fill-in across all potential numerical row pivoting during factorization.
2.5.3. Theoretical Development.
The NP-completeness of the fill-in minimization problem was established by Yannakakis (Yannakakis, 1981). Heggernes et al. (Heggernes et al., 2002) analyzed the asymptotic complexity of the minimum degree and AMD algorithms. Additionally, several theoretical developments have been made to understand the minimum degree algorithm and its variants more, including improvements to the minimum degree algorithm, randomized approaches, approximation approaches, and parameterized complexity (Cummings et al., 2021; Fahrbach et al., 2018; Natanzon et al., 1998; Bliznets et al., 2020; Cao and Sandeep, 2020; Fomin and Villanger, 2013).
2.5.4. Tie-breakings.
The importance of tie-breaking strategies in the minimum degree algorithm and its variants has long been recognized by researchers (George and Liu, 1989; Duff et al., 2017, 1976; Amestoy et al., 1996; Cavers, 1987). The quality of the resulting ordering is remarkably sensitive to the initial ordering of the input: changes of the initial ordering can increase or decrease the number of fill-ins by up to a factor of three empirically (George and Liu, 1989). Despite this sensitivity, a consistent effective tie-breaking strategy remains elusive. The state-of-the-art AMD implementation in SuiteSparse (Davis, 2019) still suffers from tie-breaking limitations. To decouple this issue when evaluating reordering algorithms, prior works (George and Liu, 1989; Amestoy et al., 1996) resorted to randomly permuting the input matrix before applying reorderings.
2.5.5. Parallelization.
Unlike nested dissection (George, 1973; Gilbert and Tarjan, 1986), another widely used fill-reducing algorithm, where extensive efforts have been made to develop parallel implementations for both shared and distributed memory systems (Chevalier and Pellegrini, 2008; Karypis and Kumar, 1998; LaSalle and Karypis, 2015), the AMD algorithm has seen little progress in this regard. To our knowledge, only two parallelization attempts related to the AMD algorithm exist. The first is in PT-Scotch (Chevalier and Pellegrini, 2008), where parallel nested dissection is followed by a variant of the AMD algorithm (Pellegrini et al., 2000) independently within each partition; however, the AMD step itself remains sequential and this paradigm relies on the existence of good separators. The second attempt is a parallelization of the column approximate minimum degree algorithm (Chen et al., 1999), in which the minimum degree criterion was removed to increase parallelism across an independent set of pivots. Nevertheless, it resulted in poor scaling and did not report the quality of the resulting orderings. Overall, to the best of our knowledge, there is currently no scalable parallel implementation of the AMD algorithm.
3. Parallel Approximate Minimum Degree Algorithm
In this section, we first explain why simply parallelizing the most computationally expensive component, approximate degree updates in Algorithm 1, is insufficient for achieveing a scalable implementation. We then describe how we address this challenge by adapting the multiple elimination strategy (Liu, 1985) to suit the approximate minimum degree (AMD) algorithm (Amestoy et al., 1996) using distance-2 independent sets. By introducing this new parallel framework and minimizing contention among threads via specialized concurrent data structures tailored to the AMD algorithm, we present the first scalable shared memory AMD implementation.
3.1. Intra-Elimination Parallelism
A straightforward approach to extract parallelism for approximate degree updates from Algorithm 1 is to parallelize by variables in line 1 and use atomics for each access to $w(e)$ between line 1 and line 3 within each elimination step. However, this strategy does not scale. Table 2 shows the average values of $|\mathcal{L}_{p}|$, $\sum_{v\in\mathcal{L}_{p}}|\mathcal{E}_{v}|$, and $\left|\bigcup_{v\in\mathcal{L}_{p}}\mathcal{E}_{v}\right|$ across all elimination steps. While the amount of parallelism, $|\mathcal{L}_{p}|$, appears reasonable, this method is not scalable since the amount of work, $\sum_{v\in\mathcal{L}_{p}}|\mathcal{E}_{v}|$, is not large enough compared to $|\mathcal{L}_{p}|$. This is also confirmed by the observation that each elimination step only takes around tens of microseconds on average. Moreover, contention is high during memory access because the number of unique elements accessed, $\left|\bigcup_{v\in\mathcal{L}_{p}}\mathcal{E}_{v}\right|$, is significantly smaller than the amount of parallelism, $|\mathcal{L}_{p}|$. These factors together illustrate why simply parallelizing the approximate degree updates fails to deliver a scalable implementation.
Table 2. Average sizes of the sets across all elimination steps, representing the amount of parallelism, the amount of work, and the number of unique elements accessed, respectively, by applying intra-elimination parallelism on Algorithm 1.
Matrix Name $|\mathcal{L}_{p}|$ $\sum_{v\in\mathcal{L}_{p}}|\mathcal{E}_{v}|$ $\left|\bigcup_{v\in\mathcal{L}_{p}}\mathcal{E}_{v}\right|$
nd24k 329.7 587.5 14.0
Flan_1565 43.8 64.8 10.2
nlpkkt240 80.5 542.8 56.3
3.2. Inter-Elimination Parallelism via Distance-2 Independent Sets
Since extracting parallelism within a single elimination step has proven unscalable, we now turn to the possibility of parallelizing the AMD algorithm across elimination steps using multiple elimination (Liu, 1985) discussed in Section 2.3. While the multiple minimum degree algorithm (Liu, 1985) improved sequential performance through multiple elimination, the AMD algorithm did not incorporate this technique. This is primarily because the main advantage of multiple elimination, reducing the cost of degree updates, diminishes when degrees are already approximated (Amestoy et al., 1996). However, multiple elimination provides an additional benefit: it enables parallelism across elimination steps by breaking dependencies between pivot eliminations.
To effectively leverage multiple elimination for parallelizing the AMD algorithm, we need to adapt the policy for selecting the set of pivots. In the original design, performance gains from multiple elimination stemmed from the overlap of pivots’ neighborhoods as explained in Section 2.3. However, such overlap is not well-suited to the AMD algorithm, as the degree of a variable is intended to be approximated with respect to the single adjacent pivot. Moreover, overlapping neighborhoods introduce significant contention during connection updates, since a variable’s updates can be affected by multiple pivots, even when those pivots do not directly interfere with each other. To address these issues, we instead select a distance-2 independent set of pivots, ensuring no overlap in the pivots’ neighborhoods as shown in Figure 4. Pivots in the distance-2 independent sets are then distributed among threads and processed independently in parallel.
Figure 4. When multiple elimination is applied to an independent set $\{1,3,7,9\}$, memory contention arises because multiple pivots attempt to update the connections of shared neighbors—variables 2, 4, 6, and 8. Moreover, since these variables are adjacent to more than one pivot, computing their approximate degrees becomes cumbersome. In contrast, using multiple elimination on a distance-2 independent set $\{1,9\}$ avoids these issues entirely, where variables 2 and 4 are only adjacent to pivot 1 and variables 6 and 8 are only adjacent to pivot 9.
[…]
An example to demonstrate the difference when multiple elimination is applied to a distance-2 independent set instead of an independent set.
We also relax the minimum approximate degree criterion for pivot selection by a multiplicative factor $mult$. Specifically, we select a distance-2 independent set from the candidate variables whose approximate degrees are no greater than $mult$ times the current minimum approximate degree. Table 3 shows the average sizes of maximal distance-2 independent sets across elimination steps for various values of $mult$, where all variables in the set are eliminated in each step. The results show that relaxation is necessary to yield enough pivots to fully utilize all threads when pivots are processed in parallel. However, overly aggressive relaxation can significantly degrade the ordering quality. We will further analyze the trade-off between relaxation factor, ordering quality, and scalability in Section 4.5.
Table 3. Average sizes of maximal distance-2 independent sets across all elimination steps with various values of multiplicative relaxation factor, $mult$, where all variables in the set are eliminated in each step.
Matrix Name $mult=1.0$ $mult=1.1$ $mult=1.2$
nd24k 2.2 9.0 10.9
Flan_1565 42.0 448.5 678.1
nlpkkt240 57.5 4084.5 6695.8
3.3. Parallelizing the Core of the Approximate Minimum Degree Algorithm
Once multiple elimination via distance-2 independent sets is applied, most components of the AMD algorithm, including approximate degree updates and the other techniques mentioned in Section 2.4, can be embarrassingly parallelized in shared memory where each thread eliminates a pivot at a time until the distance-2 independent set is depleted. However, two components still suffer from significant contention and dynamic memory management: connection updates and the maintenance of approximate degree lists. In the following, we explain how we overcome these challenges by leveraging specific properties of the AMD algorithm.
3.3.1. Concurrent Connection Updates
In the sequential AMD implementation from SuiteSparse (Davis, 2019), the sets $\mathcal{A},\mathcal{E}$ and $\mathcal{L}$ in the quotient graph are stored in a single compressed sparse column (CSC) structure with additional splitters to track the sizes of $\mathcal{A}_{v}$, $\mathcal{E}_{v}$ and $\mathcal{L}_{e}$ for each variable $v$ and element $e$. As these adjacency sets can grow beyond their initial sizes during connection updates, extra space—referred to as elbow room—is pre-allocated at the end of the CSC structure. When an adjacency set expands, it acquires space from the elbow room and migrates its contents, leaving the old space unused. Once the elbow room is exhausted, garbage collection is triggered to compact the CSC structure and reclaim unused memory, emptying elbow room at the end. Thanks to George and Liu’s result (George and Liu, 1981), which showed that the quotient graph requires no more storage than the original graph, this memory reuse mechanism ensures that no additional memory allocation is needed.
While it is possible to parallelize this structure using atomics to coordinate thread access to the elbow room, the approach suffers during garbage collection, as threads must synchronize and redo their ongoing updates after garbage collection is complete. Instead, we observe empirically that the total extra memory required during elimination never exceeds 1.5 times the storage size of $E^{0}$. This is primarily due to two facts: (i) for all variables $v$, the combined size $|\mathcal{A}_{v}|+|\mathcal{E}_{v}|$ never increases during elimination, and extra memory is only needed when creating $\mathcal{L}_{p}$ for a pivot $p$, whose size does not grow thereafter; and (ii) variables can be merged during elimination using techniques mentioned in Section 2.4, which reduces the number of pivots and, consequently, the amount of extra memory required. Therefore, by augmenting the CSC structure by 1.5 times in advance, we can empirically eliminate the need for garbage collection. To further reduce contention, each thread performs only a single atomic operation to claim extra space after collecting all connection updates associated with the assigned pivots from the distance-2 independent set. Notably, the use of a distance-2 independent set ensures that each variable is associated with at most one pivot, thereby minimizing contention.
Algorithm 2 Concurrent Approximate Degree Lists
[…]
$affinity(v)\leftarrow-1$EndFunction
[…]
3:$dlist_{tid}(loc(v)).remove(v)$EndIf
4:$dlist_{tid}(deg).insert(v)$
5:$loc(v)\leftarrow deg$
6:$affinity(v)\leftarrow tid$EndFunction
[…]
8:$dlist_{tid}(deg).remove(v)$
[…]
10:return entries in $dlist_{tid}(deg)$EndFunction
[…]
Algorithm 3 Parallel Distance-2 Independent Set with known approximate minimum degree $amd$, an analog of Luby’s algorithm (Luby, 1985)
[…]
2:     $Candidates\leftarrow[]$Forinteger $d\in[amd,\lfloor mult\times amd\rfloor$]
[…]
4:     $Candidates.resize(lim)$
[…]
6:     $l(v)\leftarrow(rand(),v)$
7:     $\forall u\in\{v\}\cup N_{v},l_{min}(u)\leftarrow(\infty,\infty)$EndFor
[…]
9:     $\forall u\in\{v\}\cup N_{v}$, atomically update $l_{min}(u)$
10:      by the minimum of $l_{min}(u)$ and $l(v)$EndFor
[…]
3.3.2. Concurrent Approximate Degree Lists
In the sequential AMD implementation from SuiteSparse (Davis, 2019), pivot selection is facilitated by maintaining $n$ degree lists. Each list corresponds to a specific approximate degree and stores the associated variables using a doubly linked list. When a pivot is selected, it is removed from the degree list, and for each variable in its neighborhood, it is moved from its current list to the new one based on its updated approximate degree. No dynamic memory allocation is required, as the total number of entries across all lists is bounded by $n$. However, this data structure is not amenable to efficient parallelization due to high contention when multiple threads attempt to access the same doubly linked list.
Instead, we let each thread $tid$ maintain its own approximate degree lists $dlist_{tid}(0,\cdots,n-1)$, along with a shared $affinity$ array to track which thread holds the most up-to-date information for each variable, and a private $loc$ array to record local degree list location of each variable. Both the $affinity$ and $loc$ arrays are initialized to $-1$. The use of a distance-2 independent set of pivots ensures that each variable is associated with at most one pivot—and thus with a single thread—during any given elimination step. As shown in Algorithm 2, when a thread attempts to remove a variable $v$, it sets $affinity(v)$ to $-1$, effectively invalidating all entries of $v$ across all threads without physically modifying the lists. When inserting $v$, the thread first checks if $v$ is currently used in its own lists based on $loc(v)$. If so, the stale entry of $v$ is explicitly removed to reclaim memory. In all cases, the variable $v$ is inserted into the appropriate list for the thread, with $loc(v)$ updated accordingly and $affinity(v)$ set to $tid$ to reflect the change. Other invalid entries, detected via mismatched affinity, are lazily reclaimed during list traversal for pivot selection. This design avoids contention for degree list maintenance, with the only necessary coordination occurring when determining the minimum approximate degree from the per-thread lists at the start of each elimination step.
3.4. Parallel Distance-2 Independent Set
To parallelize the selection of a distance-2 independent set for multiple elimination, we first collect candidate variables whose approximate degrees fall within the relaxed threshold described in Section 3.2. This collection is performed in parallel from each thread’s degree lists. To limit overhead compared to the sequential AMD algorithm, we restrict each thread to collect at most $lim$ candidates. The impact of this parameter will be analyzed in Section 4.5.
To select a distance-2 independent set, we perform a single iteration of the distance-2 analog of Luby’s algorithm (Luby, 1985) as shown in Algorithm 3. For each candidate gathered from the approximate degree lists, a random number is generated and compared against the candidates within its distance-2 neighborhood, with ties broken by candidate indices. A candidate is considered valid if and only if it has the smallest value among the candidates within its distance-2 neighborhood. To enable parallelism, Algorithm 3 includes appropriate atomic operations and synchronization barriers. We intentionally avoid running the full algorithm because a maximal distance-2 independent set is unnecessary; as long as the set obtained in the first iteration is sufficiently large, it provides enough parallelism for the core AMD algorithm.
3.5. Overall Parallel AMD Algorithm
Algorithm 4 summarizes our strategies used to parallelize the AMD algorithm. These techniques collectively enable an efficient and scalable implementation, addressing long-standing challenges in parallelizing the AMD algorithm.
Algorithm 4 Parallel Approximate Minimum Degree Algorithm
$|V|>0$
[…]
3:     Eliminate $p$ using the core AMD algorithm
4:      with concurrent connection updates and
5:      concurrent approximate degree lists
6:      (see Section 3.3)
[…]
4. Experiments
4.1. Evaluation Platforms
We evaluated the performance of our parallel approximate minimum degree algorithm on a CPU node of NERSC Perlmutter, an HPE Cray EX supercomputer. One CPU node has 128 physical cores divided between two AMD EPYC 7763 (Milan) CPUs with 512 GB of DDR4 memory in total. We used a single CPU with 64 cores in our evaluation. Our implementation was written in C++ using OpenMP for shared memory parallelism and compiled with GCC 13.2.1 using the flags -O3 -fopenmp, and it is available online at https://github.com/PASSIONLab/ParAMD. As a baseline, we used the sequential implementation from SuiteSparse (v7.10.1) (Davis, 2019). To the best of our knowledge, no existing parallel implementation of the approximate minimum degree algorithm (AMD) is available for comparison.
4.2. Matrix Suite
The sparse matrices used in our experiments were selected from the SuiteSparse Matrix Collection (Davis and Hu, 2011) to represent a diverse range of large-scale linear systems arising from real-world applications. These matrices are summarized in Table 4. Although some matrices are nonsymmetric, the AMD algorithm can still be applied by considering the sum of the matrix and its transpose, as is done in the SuiteSparse AMD implementation. In fact, this pre-processing step is performed in SuiteSparse AMD regardless of whether the input matrix is symmetric or not. For fairness in end-to-end comparison, we also parallelize this pre-processing step using simple atomic operations and include its runtime in the reported results.
Table 4. Matrices used in our experiments sorted by #nonzeros. Matrices pre2, HV15R, and stokes are nonsymmetric.
Matrix Name #rows #nonzeros Problem Type
pre2 659K 5.83M circuit simulation
nd24k 72.0K 28.7M 3D mesh
ldoor 952K 42.5M structural
dielFilterV3real 1.10M 89.3M electromagnetics
Flan_1565 1.56M 114M structural
HV15R 2.02M 283M fluid dynamics
Queen_4147 4.15M 317M structural
stokes 11.4M 349M semiconductor
nlpkkt240 28.0M 761M optimization
4.3. End-to-End Comparison
We now present the performance of our parallel AMD implementation compared to the SuiteSparse baseline, as summarized in Table 5. Matrices were randomly permuted five times with all methods evaluated on the same set of permutations to decouple tie-breaking issues explained in Section 2.5.4. We used the default relaxation factor $mult=1.1$ and set the limitation factor $lim$ to $8192$ divided by the number of threads, following the heuristic that this configuration targets a fill-in ratio of $\times1$ over the SuiteSparse baseline while ensuring a sufficiently large pool of candidates is considered for the distance-2 independent set in each elimination step. Note that with this choice of $lim$, the number of candidates remains nearly constant, regardless of the number of threads used.
Table 5. End-to-end comparison between SuiteSparse AMD and our parallel AMD implementation with 1 thread (1t) and 64 threads (64t). Matrices were randomly permuted five times with all methods evaluated on the same set of permutations to decouple tie-breaking issues. We report the median ordering time and fill-in across the five runs. Speedups and fill-in ratios are computed directly based on the medians.
Matrix Name Ordering Time (sec) Our Speedup #Fill-ins Our Fill-in Ratio
SuiteSparse Ours over SuiteSparse SuiteSparse Ours over SuiteSparse
1t 64t 1t 64t 1t 64t 1t 64t
pre2 2.01 4.70 0.62 0.43$\times$ 3.25$\times$ 1.84e+08 1.93e+08 1.91e+08 1.05$\times$ 1.04$\times$
nd24k 0.82 1.01 0.26 0.81$\times$ 3.15$\times$ 5.09e+08 5.23e+08 5.07e+08 1.03$\times$ 1.00$\times$
ldoor 1.37 1.35 0.29 1.01$\times$ 4.79$\times$ 1.52e+08 1.57e+08 1.58e+08 1.03$\times$ 1.04$\times$
dielFilterV3real 2.63 2.78 0.46 0.95$\times$ 5.74$\times$ 9.65e+08 1.06e+09 1.07e+09 1.10$\times$ 1.11$\times$
Flan_1565 4.20 4.51 0.80 0.93$\times$ 5.28$\times$ 3.72e+09 3.96e+09 4.01e+09 1.07$\times$ 1.08$\times$
HV15R 11.75 13.86 2.15 0.85$\times$ 5.46$\times$ 5.65e+10 5.65e+10 5.80e+10 1.00$\times$ 1.03$\times$
Queen_4147 13.82 17.16 2.46 0.81$\times$ 5.61$\times$ 7.72e+10 8.35e+10 8.63e+10 1.08$\times$ 1.12$\times$
stokes 79.32 121.06 10.63 0.66$\times$ 7.46$\times$ 7.76e+10 8.23e+10 8.00e+10 1.06$\times$ 1.03$\times$
nlpkkt240 296.67 433.95 35.75 0.68$\times$ 8.30$\times$ 1.60e+12 1.83e+12 1.79e+12 1.14$\times$ 1.12$\times$
Table 5 shows that, when running with a single thread, our implementation incurs overhead, resulting in slower ordering time. This is primarily due to the inclusion of Algorithm 3 for selecting distance-2 independent sets, a step not required in the original AMD algorithm. However, it is this modification that enables effective parallelization, achieving up to an 8.30×8.30$\times$8.30 × speedup over SuiteSparse and up to a 12.21×12.21$\times$12.21 × self-speedup when using 64 threads. Our fill-in ratio is consistently around $\times1$ over SuiteSparse, reflecting the effect of the relaxation factor $mult$ in avoiding the selection of overly suboptimal pivots.
4.4. Runtime Breakdown and Parallelism Analysis
To gain deeper insights into the performance of our parallel AMD implementation, we present a runtime breakdown in Figure 5 using the default parameters.
First, we observe that the pre-processing step, which computes $A+A^{T}$, becomes a bottleneck for some matrices and does not scale well with increasing thread count. This behavior is expected due to high contention and irregular memory access patterns inherent to the operation. However, this step is unnecessary for symmetric matrices and, for nonsymmetric matrices including pre2, HV15R, and stokes, its impact is minimal. We include it in the runtime solely for a fair end-to-end comparison with the SuiteSparse baseline.
Second, while the core of the AMD algorithm is highly parallelizable in theory, it does not achieve perfect scaling in practice. For example, on the nd24k matrix, we observe only a 3.7×3.7$\times$3.7 × speedup when using 64 threads for the core AMD part. This limited scalability is explained by the distribution of the sizes of the distance-2 independent sets, as shown in Figure 6, which directly affects the amount of parallel work available in each elimination step. Unsurprisingly, the distance-2 independent sets selected for nd24k are relatively small compared to other matrices, resulting in underutilization of threads. Moreover, across all matrices, a significant portion of the distance-2 independent sets still have sizes less than 64, further illustrating why the core AMD algorithm does not scale perfectly despite being embarrassingly parallel. As a future direction, we propose dynamically adapting the relaxation and limitation factors during elimination, particularly when low workload is detected, to extract greater parallelism when needed.
Figure 5. Runtime breakdown of our parallel AMD algorithm as the number of threads scales from 1 to 64.
[…]
Runtime breakdown of our parallel AMD algorithm as the number of threads scales from 1 to 64.
Figure 6. Violin plots showing the distribution of the sizes of distance-2 independent sets across elimination steps. The plots are cropped to reflect the actual range of the data. The dotted line marks the threshold of 64, representing the minimum size needed to fully utilize all 64 threads.
[…]
Violin plots showing the distribution of distance-2 independent set sizes across elimination steps. The dotted line marks the threshold of 64, representing the minimum size needed to fully utilize all 64 threads.
Figure 7. Impact of the relaxation factor $mult$ and the limitation factor $lim$ on core AMD runtime, distance-2 independent set selection time, and ordering quality. Results are shown for matrices nd24k and nlpkkt240, representing the worst and best scalability cases, respectively. All experiments were conducted using 64 threads. Lighter colors indicate better performance.
[…]
Impact of the relaxation factor $mult$ and limitation factor $lim$ on core AMD runtime, distance-2 independent set selection time, and ordering quality. Results are shown for matrices nd24k and nlpkkt240, representing the worst and best scalability cases, respectively. All experiments were conducted using 64 threads. Lighter colors indicate better performance.
4.5. Relaxation and Limitation Factors
We now evaluate the impact of the relaxation factor $mult$ and the limitation factor $lim$. Figure 7 illustrates how these parameters affect the runtime of the core AMD algorithm, the runtime of distance-2 independent set selection, and the overall ordering quality. Our analysis focuses on two representative matrices using 64 threads: nd24k, which demonstrates the worst scalability, and nlpkkt240, which shows the best scalability, in Table 5.
The impact of the relaxation factor $mult$ is more pronounced and consistent. When $mult$ is too small, the algorithm lacks sufficient parallelism, leading to poor performance. On the other hand, excessively large values of $mult$ degrade the ordering quality due to the inclusion of suboptimal pivots.
In contrast, the impact of the limitation factor $lim$ is less straightforward. Its effects are more nuanced and do not follow a universal trend. This is because adjusting $lim$ affects several aspects of the algorithm, including but not limited to the total number of elimination steps, opportunities to apply other orthogonal techniques discussed in Section 2.4, and the amount of computation available to hide memory access latency. As these factors impact different matrices to varying degrees, the limitation factor $lim$ produces complex behavior with no single dominant pattern.
Nevertheless, Figure 7 suggests that the optimal configuration occurs around $mult=1.2$ and $lim=128$. After further fine-grained tuning across our dataset, we found that setting $mult=1.1$ provides a better balance between runtime and ordering quality. To ensure that the number of candidates for the distance-2 independent set remains consistent regardless of the number of threads, we finalize $lim$ as 8192 divided by the number of threads. This results in $lim=128$ when using 64 threads, aligning with the identified optimal value. These together result in the default settings we used for the other experiments.
5. Conclusion and Future Work
We have presented the first scalable shared memory AMD implementation. By incorporating distance-2 independent sets into multiple elimination and designing specialized concurrent data structures tailored to the AMD algorithm, we have enabled effective parallelism while reducing contention. Our experiments have demonstrated that this approach scales and consistently outperforms the long-standing AMD implementation from SuiteSparse, with only modest and controllable overhead in fill-in.
As immediate future work, we plan to further investigate the impact of the limitation factor and explore strategies for dynamically adapting both the relaxation and limitation factors during elimination, particularly when low workload is detected, to extract greater parallelism when needed. In the longer term, we aim to address tie-breaking issues and develop an efficient, parallel-friendly approach for this critical component.
Acknowledgments
This research is supported by the Applied Mathematics program of the Advanced Scientific Computing Research (ASCR) within the Office of Science of the DOE under Award Number DE-AC02-05CH11231. We used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award ASCR-ERCAP-33069.